# Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors

### v1

@@ 3.1.2 Optimization Using Wake-Sleep in Latent Space
This procedure mirrors wake-sleep; however, sampling $\mathbf{x}$ from $p_{\bm{\theta}}$ to obtain gradients for the sleep term introduces computational overhead. To address this issue, we propose wake-sleep in latent space, an algorithm that optimizes an approximation ^$\hat{\mathcal{L}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ of $\mathcal{L}$:
@@
^$\hat{\mathcal{L}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ $=\underbrace{\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / reconstr. term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}-\underbrace{D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}-\underbrace{D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))}_{\text{latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$}}.$ (7)
@@
3.1.3 DDVI with Diffusion-Based Encoders
Our framework is naturally instantiated with diffusion models. Let $\mathbf{y}_{0}=\mathbf{z}$ and $\mathbf{y}_{T}=\mathbf{y}$. The forward noising model can be defined as $r(\mathbf{y}_{1:T}|\mathbf{z})=\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1})$, where the $\mathbf{y}_{0:T}$ are an extended set of latent variables that represent increasingly noised versions of $\mathbf{y}_{0}$. We parameterize the approximate reverse diffusion process as $q_{\bm{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})=q(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x}).$
[…]
$\mathcal{L}_{\text{diff}}=$ $\mathbb{E}_{r}\left[\log q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})-\sum_{t=2}^{T}D_{\mathrm{KL}}(r(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{y}_{0})||q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x}))\right]-D_{\mathrm{KL}}(r(\mathbf{y}_{T}|\mathbf{z})||q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})).$ (10)
This bound is an instantiation of Equation 8 when $r$ is a diffusion process. Many choices of $r,q$ fit this framework. A common type of noising process is Gaussian diffusion, where we define $r(\mathbf{y}_{t}|\mathbf{y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{1-{\bm{\alpha}}_{t}}\mathbf{y}_{t-1},{\bm{\alpha}}_{t}\mathbf{I})$ for a suitable schedule $({\bm{\alpha}}_{t})_{t=1}^{T}$. We then adopt the parameterization $q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})=\mathcal{N}(\mathbf{y}_{t-1};\mu_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x},t),\Sigma_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x},t))$. It is then common to parameterize $q_{\bm{\phi}}$ with a noise prediction network ${\bm{\epsilon}}_{\bm{\phi}}$ (Ho et al., 2020); the sum of KL divergences can be approximated by $E_{t,{\bm{\epsilon}}_{t}\sim r(\mathbf{y}_{0},t)}||{\bm{\epsilon}}_{t}-{\bm{\epsilon}}_{\bm{\phi}}(\sqrt{\bar{{\bm{\alpha}}_{t}}}\mathbf{y}_{0}+\sqrt{1-\bar{{\bm{\alpha}}_{t}}}{\bm{\epsilon}}_{t},\mathbf{x},t)||^{2}$. By our earlier argument, this objective encourages $q_{\bm{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})$ to match $r(\mathbf{y}_{0:T})$ as well as the true posterior $p(\mathbf{z}|\mathbf{x})$.
Lastly, we need to show that this choice of $q$ admits a tractable entropy. This follows from
$H(q)$ $=-\sum_{t=1}^{T+1}\mathbb{E}_{q}[\log q(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})]=\sum_{t=1}^{T+1}\mathbb{E}_{q}[\frac{d}{2}\left(1+\log(2\pi)\right)+\frac{1}{2}\log|\Sigma_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x})|]$ (11)
3.2 Auxiliary-Variable Diffusion Encoders
Denoising diffusion variational inference can also be understood from the perspective of auxiliary-variable generative models (Maaløe et al., 2016). Our method can be seen as introducing an approximate posterior $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})=\int_{\mathbf{y}}q_{\bm{\phi}}(\mathbf{z},\mathbf{y}|\mathbf{x})$ which itself contains latent variables $\mathbf{y}$.
The standard approach to fit auxiliary-variable generative models is to apply the ELBO twice:
$\log p_{\bm{\theta}}(\mathbf{x})$ $\geq\log p_{\bm{\theta}}(\mathbf{x})-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))=\text{ELBO}_{\mathbf{z}|\mathbf{x}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ (12)
$\geq\text{ELBO}_{\mathbf{z}|\mathbf{x}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})-\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$ (13)
$=\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x},\mathbf{z})r(\mathbf{y}|\mathbf{x},\mathbf{z})-\log q(\mathbf{y},\mathbf{z}|\mathbf{x})],$ (14)
where in Equation 13 we applied the ELBO over $\mathbf{z}$, and in Equation 12 we applied the ELBO again over the latent $\mathbf{y}$ of $q$. The $r(\mathbf{y}|\mathbf{x},\mathbf{z})$ can be interepreted as an approximate variational posterior for the true posterior $q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})$ of the data distribution. Specifically, the gap between Equation 12 and $\log p_{\bm{\theta}}(\mathbf{x})$ is precisely $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))+\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$. Thus, if we correctly match the pairs of approximate posteriors, we will achieve a tight bound.
Additionally, notice that Equation 14 is exactly the ELBO in Equation 2 after rearranging terms. Thus the two perspectives are equivalent: using variational inference to fit an augmented model $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})$ is equivalent to fitting the original $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ and applying the ELBO to $p_{\bm{\theta}}$, followed by applying the ELBO again to the latent $\mathbf{y}$ in $q$. Since our objective $\mathcal{L}$ is Equation 14) augmented with $\mathcal{L}_{\text{sleep}}$, denoising VI can also be interpreted as fitting an latent-variable approximate posterior $q$. As seen in Equation 9, we implicitly minimize the reverse KL divergence $\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})}[D_{\mathrm{KL}}(r(\mathbf{y}|\mathbf{z})||q_{\bm{\phi}}(\mathbf{y}|\mathbf{z},\mathbf{x}))]$.
Note, however, that there is a crucial difference between our method and auxiliary-variable generative models: we do not optimize the parameters of $r$. Rather, $r$ is a user-defined noising process that acts as a regularizer for the model $q_{\bm{\phi}}$ and whose reverse we seek to fit.
3.3 Regularizing Variational Inference with Denoising Processes
Lastly, our approach can be understood from a third perspective: that of regularization. As we mentioned earlier, $\mathcal{L}$ is the ELBO with an additional regularization term that encourages the approximate posterior $q(\mathbf{z}|\mathbf{x},\mathbf{y})q(\mathbf{y}|\mathbf{x})$ to gradually transform $\mathbf{y}$ into $\mathbf{z}$ following the guidance of $r$. In that sense, $r$ provides data augmentation to ensure that the mapping from $\mathbf{y}$ to $\mathbf{x}$ is easier to learn, similarly to how a diffusion process provides additional learning signal for generating $\mathbf{x}$ from noise.
Our framework also supports the user defining general distributions as regularizers $r_{1}(\mathbf{z},\mathbf{y})$ and $r_{2}(\mathbf{y}|\mathbf{z},\mathbf{x})$. Here, $r_{1}$ puts a custom prior on $\mathbf{z},\mathbf{y}$, while $r_{2}$ specifies a mapping between $\mathbf{y},\mathbf{z}$ that should be followed by the encoder model. We write $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}},r_{1})$ and $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}},r_{2})$, where each term is still defined as in Equation 3 and Equation 6, except using $r_{1},r_{2}$ and we optimize
$\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}},r_{1},r_{2})=\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})+\beta\cdot\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}},r_{1})+\gamma\cdot\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}},r_{2}),$ (15)
where $\beta,\gamma>0$ (in particular, $\beta$ mirrors the $\beta$-VAE model (Higgins et al., 2016)).
Examples of $r_{2}$ include Gaussian diffusion, discrete noising processes, as well as noise conditioned on $\mathbf{x}$. Note that $r_{2}$ could even be fully deterministic when parameterized with a bijective flow. With a strong regularizer, we could set $q(\mathbf{z}|\mathbf{y},\mathbf{x})$ to equal the reverse $r_{2}(\mathbf{z}|\mathbf{y},\mathbf{x})$ of this flow. Thus, for suitably designed $r_{2}$, we could perfectly match the shape of any mapping. This mirrors the InteL-VAE framework (Miao et al., 2021), with support for flexible and probabilistic regularizers.
Most importantly, our method can be understood as introducing a new form of regularization. While a standard VAE introduces a prior regularization term, we introduce a new form of regularization via a function $r_{2}$. This can be advantageous when specifying or enforcing certain priors is difficult.
4 Extensions
4.1 Semi-Supervised Learning
Following Makhzani et al. (2015), we extend our algorithm to the semi-supervised learning setting where some data points are labeled and we define $p(\mathbf{z},l)=p(\mathbf{z}|c)p(l)$. Then, the model can be specified as $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},l)=p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},l)r(\mathbf{y}|\mathbf{z},l)p_{\bm{\theta}}(\mathbf{z}|l)p(l)$ and the variational distributions are $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y},l),q_{\bm{\phi}}(\mathbf{y}|\mathbf{x}),q_{\bm{\phi}}(l|\mathbf{x})$. In this setting, we consider two cases of whether the label can be observed or not (Kingma et al., 2014). We extend Equation 7 to incorporate the label $l$ corresponding to a data point as follows:
$\mathcal{L}_{\text{semi}}=\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},l)]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}|l))-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|l)||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},l))$
4.2 Clustering
We have further extended our algorithm to encompass the clustering paradigm. We propose two distinct strategies. In the first approach, we simply set $p_{\bm{{\bm{\theta}}}}(\mathbf{z})$ as a mixture of desired priors. The means of these priors are characterized by ${\bm{{\bm{\theta}}}}$. From these means, cluster membership, denoted as $\mathbf{c}$ can be deduced. Remarkably, this approach requires no alteration to the existing learning objective.
$\mathcal{L}_{\text{clus}}(\mathbf{x})$ $=\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z},\mathbf{c}))-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))$
5 Experiments
We compare DD-VAE with vanilla VAE (Kingma & Welling, 2013), IAF-VAE (Kingma et al., 2016), and AAE (Makhzani et al., 2015) on MNIST (Lecun et al., 1998) and CIFAR-10 (Krizhevsky & Hinton, 2009) in unsupervised and semi-supervised learning settings, and also on the 1000 genomes dataset Siva (2008) in clustering settings. The differences between these methods are summarized in Table 8. We discuss the computational costs of all methods in Section A.1. The priors, model architecture, and training details can be founded in Section A.2, Section A.3, and Section A.4 respectively. All results below are reported with 95% confidence interval using 5 different seeds.
5.1 Unsupervised learning
We fit a model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ on the MNIST and CIFAR-10 datasets with three priors $p(\mathbf{z})$: pinwheel, swiss roll, and square and report our results in Table 1 and Table 5. We measure a k-nearest neighbors classification accuracy of the latents (Acc). We also measure latent negative log-likelihood (Latent NLL) by fitting a kernel density estimator (KDE) on the latents produced by the model with test data as input and compute the log-likelihood of the latents sampled from the prior under the fitted KDE.
We consistently see our method perform best on both KNN accuracy and latent NLL, indicating that its output latents are the most informative of labels and best-aligned with the priors. Table 1 reports qualiatitve results that indicate that DD-VAE enforces most accurately the correct prior shape.
Figure 2: Unsupervised visualization on MNIST using three different priors (pinwhell, swiss roll, and square). Each color indicates a class.
Method Pinwheel Swiss Roll Square Acc Latent NLL Acc Latent NLL Acc Latent NLL VAE $0.58\pm 0.01$ $6.18\pm 0.93$ $0.56\pm 0.055$ $82.96\pm 99.30$ $0.35\pm 0.02$ $25.07\pm 62.41$ IAF-VAE $\bm{0.71\pm 0.00}$ $2.06\pm 0.21$ $0.67\pm 0.02$ $12.13\pm 18.28$ $0.65\pm 0.03$ $1.53\pm 0.82$ AAE $0.67\pm 0.02$ $1.94\pm 0.09$ $0.68\pm 0.03$ $3.43\pm 0.14$ $0.57\pm 0.02$ $1.97\pm 0.70$ HVAE w/ flow $\bm{0.72\pm 0.01}$ $2.18\pm 0.13$ $\bm{0.71\pm 0.02}$ $4.17\pm 1.23$ $\bm{0.70\pm 0.01}$ $\bm{1.25\pm 0.08}$ DiffVAE $\bm{0.72\pm 0.01}$ $\bm{1.57\pm 0.10}$ $0.66\pm 0.01$ $\bm{3.19\pm 0.07}$ $\bm{0.72\pm 0.03}$ $\bm{1.24\pm 0.13}$
Table 1: Unsupervised learning on MNIST. We report accuracy using KNN (K=20) classifier and latent negative log-likelihood with pinwheel, swiss roll, and square priors.
5.2 Semi-supervised Learning
We also evaluate the performance of our method and the baselines under semi-supervised learning setting where some labels are observed (1,000 for MNIST and 10,000 for CIFAR-10) and the partitions of the priors are known. We use the same set of priors and baselines. Details on how we partition each prior into $p_{\bm{{\bm{\theta}}}}(\mathbf{z}|l)$ can be founded in Section A.2. The partitions defined for our priors are local parts of the priors. We substitute the KNN accuracy metric with the average class latent NLL, which measures how well the test latents from each class match their respective partition of the prior $p_{\bm{{\bm{\theta}}}}(\mathbf{z}|l)$. In our case, this metric implicitly measures accuracy because the metric requires that the test latents for each class be close to each other, as the partitions defined are local parts of the full prior. The results are shown in Table 2 and Table 6. DD-VAE mostly outperforms the baselines across different priors and metrics. We also show the visualization in Figure 3 where DD-VAE matches the prior almost perfectly.
Method Pinwheel Swiss Roll Square Latent NLL Avg Class Latent NLL Avg Class Latent NLL Avg Class Latent NLL Latent NLL Latent NLL VAE $7.77\pm 1.83$ $15.64\pm 8.79$ $56.91\pm 33.90$ $244.51\pm 94.19$ $1.30\pm 0.03$ $-0.71\pm 0.05$ IAF-VAE $2.71\pm 0.22$ $1.84\pm 0.46$ $8.10\pm 12.17$ $14.59\pm 31.88$ $1.04\pm 0.04$ $-1.01\pm 0.03$ AAE $1.90\pm 0.04$ $0.64\pm 0.11$ $\bm{3.35\pm 0.17}$ $1.86\pm 0.77$ $1.70\pm 0.26$ $2.50\pm 4.70$ DiffVAE $\bm{1.28\pm 0.01}$ $\bm{-0.94\pm 0.01}$ $\bm{3.35\pm 0.10}$ $\bm{1.24\pm 0.09}$ $\bm{0.96\pm 0.03}$ $\bm{-1.20\pm 0.02}$
Table 2: Semi-supervised learning on MNIST (1,000 labels). We report average class latent negative log-likelihood and latent negative log-likelihood with pinwheel, swiss roll, and square priors.
5.3 Clustering and Visualization for Genotype Analysis
In this section, we report results on an real-world task in genome analysis. Visualizing genotype data reveals patterns in the latent ancestry of individuals. We compare DD-VAE against with the three strong clustering baselines using the 1000 Genomes dataset. We also report visualizations from three dimensionality reduction algorithms: PCA, TSNE, and UMAP. For each clustering algorithm, we seek to discover up to 20 clusters. We report quantitative results in terms of cluster purity, cluster completeness, and normalized mutual information (NMI). There is an inherent trade-off between cluster purity completeness. The overall clustering performance can be captured with NMI.
Method Cluster Cluster NMI
Method Purity Completeness
VAE $0.28\pm 0.02$ $\bm{0.78\pm 0.16}$ $0.59\pm 0.08$
IAF-VAE $0.29\pm 0.04$ $\bm{0.73\pm 0.06}$ $0.55\pm 0.06$
AAE $0.37\pm 0.06$ $\bm{0.76\pm 0.11}$ $\bm{0.63\pm 0.02}$
DiffVAE $\bm{0.45\pm 0.03}$ $\bm{0.75\pm 0.05}$ $\bm{0.66\pm 0.04}$
Table 3: Quantitative genotype clustering results.
In Table 3, we see that DD-VAE attains the best performance on cluster purity and NMI. For cluster completeness, VAE and AAE have better means but much larger confidence interval. Furthermore, we visualize our genotype clustering results in latent space, shown in Figure 4, and also report results from classical dimensionality reduction and visualization methods that do not perform clustering (PCA (Wold et al., 1987), t-SNE (Van der Maaten & Hinton, 2008), and UMAP (McInnes et al., 2018)). The legend of Figure 4 can be founded at Figure 5.
6 Related Work, Discussion, and Conclusion
@@
There are a number of recent works merging VAE and diffusion models. Vahdat et al. (2021); Wehenkel & Louppe (2021) use diffusion with VAE priors. Recent research (Preechakul et al., 2022; Zhang et al., 2022; Wang et al., 2023b) has also melded auto-encoders with diffusion models, focusing on semantically meaningful latents. Cohen et al. (2022) crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution. These works focus on sample quality and directly fit $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to a complex distribution (e.g., compressed images (Rombach et al., 2022)); our work instead seeks to achieve high sample quality in latent space, and obtains complex $\mathbf{z}$ by diffusion from a simple $\mathbf{y}\sim q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})$.
Latent variable models in general are an attractive alternative to visualization methods like PCA, UMAP, and t-SNE (McInnes et al., 2018; Van der Maaten & Hinton, 2008). Domain-specific knowledge can be injected through the prior, and deep neural networks can be utilized to achieve a more expressive mapping from the data space to the latent space. Nevertheless, downsides of LVMs are that they are more computationally expensive and require careful hyperparameter tuning.
While this paper focuses on applications of DD-VAE to dimensionality reduction and visualization, there exist other tasks for the algorithm, e.g., density estimation or sample quality. Since our learning objective differs from the ELBO (it adds a regularizer), we anticipate gains on models whose training benefits from regularization, but perhaps not on all models. Also, attaining competitive likelihood estimation requires architecture improvements that are orthogonal to this paper. However, our ability to generate diverse samples and achieve class separation in latent space hints at the method’s potential on these tasks. Accurate variational inference has the potential to improve downstream applications of generative modeling, e.g., decision making or causal effect estimation (Nguyen & Grover, 2022; Deshpande et al., 2022; Deshpande & Kuleshov, 2023; Rastogi et al., 2023).
Appendix A Appendix
A.1 Computational Cost Analysis

### v3

@@ Appendix B Simplifying Wake-Sleep
In wake-sleep, sampling $\mathbf{x}$ from $p_{\bm{\theta}}$ to obtain gradients for the sleep term introduces computational overhead. To address this issue, we propose wake-sleep in latent space, an algorithm that optimizes an approximation ^$\hat{\mathcal{L}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ of $\mathcal{L}$:
@@
^$\hat{\mathcal{L}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ $=\underbrace{\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / reconstr. term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))}_{\text{latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$}}.$ (16)
@@
Appendix C Comparision of Methods
We provide a comprehensive comparison of different methods in Table 3. Vahdat et al. [39], Wehenkel and Louppe [43], Rombach et al. [30] perform diffusion in the latent space of a VAE to improve the efficiency of image generation. Their goal is high sample quality, and they introduce into $p$ hierarchical latents with simple Gaussian priors. Our goal is different: we seek a method to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$.
@@ Appendix C Comparision of Methods
Recent work [25, 47, 42] has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$. Cohen et al. [5] crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution.
Model Training Objective Approximating Family Sample-based Prior Auxiliary Variable Tasks Simplified Graphical Illustration AEVB ELBO Diagonal Gaussian Density estimation $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ AEVB-IAF ELBO Normalizing flow Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{x}$ AAEB Adversarial training Adversarial generator Visualization $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ H-AEVB-(IAF) ELBO Factorial Normal / Normalizing flow Density estimation / High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ ADGM ELBO Non-Gaussian Density estimation $\mathbf{x}\rightarrow\mathbf{a}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ LDM ELBO Diagonal Gaussian High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ LSGM ELBO & score matching Diagonal Gaussian High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ DDVI ELBO & sleep term Denoising diffusion Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{T}(\mathbf{y})\rightarrow\mathbf{z}_{0}(\mathbf{z})\rightarrow\mathbf{x}$
Table 3: Comparison of DDVI to other relevant methods. $\mathbf{x}$ represents the original data input to the model. $\mathbf{z}$ denotes the latent (hidden) representation of the input data. $\mathbf{a}$ represents an auxiliary variable introduced in some models (like ADGM) to capture additional aspects of the data distribution or to assist in the model’s learning process.
Appendix D Computational Cost Analysis